OpenStudy (anonymous):

Write an equation of the line that passes through the given point and is parallel to the given line. (-2,1) , y=2x+5

OpenStudy (anonymous):

y = m x + b y = 1, m = 2 and x = -2 Plug in these values in the equation above and solve for b. It turns out that the new line is the same line, y=2x+5

OpenStudy (anonymous):

@robtobey I got a different answer. Can I show you?

OpenStudy (anonymous):

y-1=1/2(x+2) y-1=1/2x+5/2 +1 +1 y=1/2x+7/2

OpenStudy (anonymous):

The new line equation is the following. It has the same slope, is parallel to, the given line. Therefore, the slope is 2, the coefficient of x in the given line. y = m x + b y = 2 x + b To find b, use the point coordinates from the point (-2,1) and solve for b. 1 = 2*(-2) + b 1 = -4+ b 5 = b The new line is: y = 2x + 5